### v5

# Gaining Wisdom from Setbacks[Uncaptioned image]: Aligning Large Language Models via Mistake Analysis

Abstract
The rapid development of large language models (LLMs) has not only provided numerous opportunities but also presented significant challenges. This becomes particularly evident when LLMs inadvertently generate harmful or toxic content, either unintentionally or because of intentional inducement. Existing alignment methods usually direct LLMs toward the favorable outcomes by utilizing human-annotated, flawless instruction-response pairs. Conversely, this study proposes a novel alignment technique based on mistake analysis, which deliberately exposes LLMs to erroneous content to learn the reasons for mistakes and how to avoid them. In this case, mistakes are repurposed into valuable data for alignment, effectively helping to avoid the production of erroneous responses. Without external models or human annotations, our method leverages a model’s intrinsic ability to discern undesirable mistakes and improves the safety of its generated responses. Experimental results reveal that our method outperforms existing alignment approaches in enhancing model safety while maintaining the overall utility.
00footnotetext: Equal contribution: kai.chen@connect.ust.hk, wangchunwei5@huawei.com 00footnotetext: Corresponding authors: honglanqing@huawei.com, mifei2@huawei.com
1 Introduction
In recent years, large language models (LLMs) have experienced exponential growth in their capabilities, leading to significant advancements in various fields, especially in understanding and generating human-like texts (Kaddour et al., 2023; Wang et al., 2023; OpenAI, 2023). However, these achievements are also accompanied by challenges. Notably, when trained on an extensive amount of noisy web text corpora, LLMs can easily produce harmful responses even without the red-teaming prompts, posing substantial risks in downstream deployment (Parrish et al., 2021; Liang et al., 2021; Hartvigsen et al., 2022). Given the powerful capabilities of LLMs and their extensive range of applications, it becomes crucial to ensure these models operate in a manner that resonates with human morals. Aligning LLMs with human values is not merely important, it is imperative (Xu et al., 2020; Zhang et al., 2022; Dinan et al., 2022).
Existing alignment methods of LLMs mainly employ two principal methodologies: supervised fine-tuning (SFT) (Radiya-Dixit & Wang, 2020; Ouyang et al., 2022; Liu et al., 2023a) and reinforcement learning with human feedback (RLHF) (Christiano et al., 2017; Ibarz et al., 2018; Jaques et al., 2019; Bai et al., 2022a). SFT-based methods align LLMs with human values using large volumes of human-annotated instruction-response pairs (Ouyang et al., 2022), primarily teaching them to learn the nature of good responses. On the other hand, RL-based methods guide LLMs to produce appropriate responses by using reward models to select relatively better responses based on human feedback (Ibarz et al., 2018). In summary, these existing alignment methods primarily rely on large volumes of human-annotated, error-free instruction-response pairs for alignment. Harmful or erroneous data are often discarded and rarely considered for potential use in model alignment.
On the other hand, it is widely acknowledged that humans can derive profound insights from mistakes. This is echoed in an old Chinese proverb, “A fall into the pit is a gain in your wit”, emphasizing the intrinsic value of learning from mistakes for deeper understanding. However, directly exposing LLMs to erroneous instruction-response pairs using methods like SFT or RLHF may cause them to learn and replicate harmful text patterns (Liu et al., 2023a). This leads to a challenging problem: How can LLMs learn from the mistakes without being negatively influenced by toxic inputs?
Insights from human practice suggest that, learning from mistakes requires first identifying the errors, understanding their underlying causes, and then preventing them in the future. Therefore, our preliminary study begins by evaluating the LLMs’ ability to identify mistakes, which is termed “discrimination” in this work (see Sec. 3). It is compared with “generation” (i.e., a LLM’s ability to generate appropriate responses to an instruction). Specifically, we ask LLMs to analyze their own responses to red-teaming instructions related to harmful problems. Surprisingly, even models not yet aligned for safety (such as Alpaca (Taori et al., 2023)) can identify mistakes in their own responses. Details are provided in Sec. 3, which might be because “discrimination” (i.e., recognizing mistakes in responses) tends to be easier than “generation” (i.e., generating appropriate responses), thereby equipping LLMs with self-critique capabilities (Huang et al., 2022; Saunders et al., 2022; Gou et al., 2023b). Motivated by this observation, we propose a novel alignment framework that trains LLMs through self-critique and mistake analysis (see Fig. 1 as an illustration).
We start by inducing an unaligned model to generate harmful responses for mistake collection (see Sec. 5.1). Then, we inform the model about the potential mistakes and instruct it to evaluate its own responses. This mistake analysis data, along with regular helpful and harmless instruction-response pairs, is used for model fine-tuning. This allows LLMs to simultaneously learn what should and should not be generated to improve alignment performance, as the mistake analysis data serve as a “fine-grained mask” that helps avoid harmful content. Additionally, we demonstrate that our method can effectively defend post-aligned LLMs against novel instruction attacks using a limited number of representative mistakes (see Sec. 5.2). In summary, our method is the first to leverage natural-language-based mistake analysis provided by the model itself for alignment. Extensive experiments on various benchmarks (Dubois et al., 2023; Dai et al., 2023) demonstrate the superiority of our method. The main contributions of this work are threefold:
Figure 1: Pipeline of alignment with mistake analysis. Different from existing works (e.g., SFT and RLHF) that focus on aligning LLMs with “error-free responses”, the proposed method deliberately exposes LLMs to harmful contexts to learn the reasons for mistakes.
  1. 1.
    We introduce a novel alignment framework that aligns LLMs by transforming erroneous instruction-response pairs into valuable alignment data based on mistake analysis.
  2. 2.
    We are the first to demonstrate that a LLM can achieve self-alignment without external models and additional human annotations. The model’s inherent discrimination ability can be utilized to enhance its own generation capability.
  3. 3.
    Extensive experiments show that our method outperforms both SFT and RL-based methods in ensuring model safety while maintaining overall utility on various benchmarks.
2 Related Work
Supervised Fine-Tuning
(SFT) is widely used to align LLMs with human expectations (Ouyang et al., 2022; Wang et al., 2023; Gou et al., 2023a). This method involves calculating the cross-entropy loss over the ground-truth response to an input instruction, thereby training LLMs to adhere to user instructions. A key limitation of SFT is its exclusive focus on optimal responses, without comparing them with the less optimal ones. To overcome this problem, variants of SFT have been developed, such as Reward Ranked Fine-tuning (RAFT) (Dong et al., 2023) and Chain of Hindsight (CoH) (Liu et al., 2023a). RAFT employs a reward model to score and select samples, fine-tuning the model with only high-reward examples. CoH, in contrast, fine-tunes LLMs using sequences of responses along with human feedback, enabling the models to discern differences between various responses. However, these SFT-based methods primarily direct the model towards identifying an “optimal response”, and often shield it from poorer responses.
Reinforcement Learning from Human Feedback
(RLHF) (Ouyang et al., 2022) optimizes LLMs using the human-elicited reward models (RM), typically trained with pairwise human preferences for model outputs. However, obtaining high-quality human-labeled preference data at scale is still resource-intensive. RLAIF (Lee et al., 2023) addresses this by simulating human preferences using LLMs, though the labels might be noisier than human-validated ones. Further advancements in alignment include approaches like DPO (Rafailov et al., 2023) and RRHF (Yuan et al., 2023). DPO integrates ranking information into LLM fine-tuning, while RRHF modifies loss terms for improved alignment. The use of contrastive methods in reinforcement learning (Yang et al., 2023) has shown improvements in sample efficiency and model quality by highlighting the differences between positive and negative responses. While RL-based methods enable models to assess the relative quality of responses, they often do not provide specific reasons for penalizing lower-quality outputs.
Self-correction
and self-improvement are increasingly recognized capabilities of LLMs. Huang et al. (2022) demonstrate that LLMs can enhance their reasoning skills using self-generated solutions on unlabeled datasets. Gou et al. (2023b) introduce the CRITIC framework, which allows LLMs to modify their outputs through interaction with external tools, akin to human validation processes. Saunders et al. (2022) fine-tune LLMs to generate critiques, aiding human annotators in spotting content flaws. Bai et al. (2022b) confirm that LLMs could achieve moral self-correction when trained with human feedback. These findings support the notion that LLMs are capable of performing mistake analysis, offering valuable insights into their errors.
(a) (a) Comparison between generation and discrimination abilities for Alpaca, GPT-3 and GPT-3.5. (b) Comparison between guided and unguided analyses. Check details in Sec. 3.
3 Preliminary
3.1 Generation against Discrimination
In this section, we examine a LLM’s mistake detection capabilities, which is termed “discrimination” in our work. It is compared with “generation”, denoting the LLM’s ability to generate appropriate responses to a given instruction. Specifically, in our work, “discrimination” involves more than just categorizing a response as “harmful” or “harmless”. It requires a nuanced analysis of the response quality, as demonstrated in Fig. 7(b). Our focus is on comparing a LLM’s ability to generate appropriate responses versus its ability to identify potential mistakes.
Three models are considered, including Alpaca (Taori et al., 2023), GPT-3 (Olmo et al., 2021) and GPT-3.5 (Ye et al., 2023). These models are evaluated with 500 red-teaming instructions from the PKU-SafeRLHF dataset, which usually involves harmful problems. The models then analyze the safety of their own responses. We compare the quality between two types of pairs: (1) Instruction-Response pairs for generation, and (2) (Instruction, Response)-Analysis pairs for discrimination. For evaluation, GPT-4111https://chatgpt.ust.hk is used to rate the quality of these pairs on a scale of 1 to 10. This is followed by a human verification process. For detailed experimental settings, please refer to Appendix A.1.
As shown in Fig. LABEL:fig:pre_gen_vs_dis, across all the evaluated models, the discrimination scores (i.e., identifying and analyzing potential mistakes) consistently exceed those of generation (i.e., producing harmless responses directly) with a significant margin. This indicates that even though LLMs may generate harmful responses, they still have the capability to identify the harmful elements within their own responses (see examples in Appendix A.1). This observation supports the idea that discrimination is simpler than generation (Saunders et al., 2022).
3.2 Guided Analysis against Unguided Analysis
We further explore how to boost the inherent discrimination ability of LLMs. We evaluate a LLM’s capability to analyze potential mistakes in two different scenarios: (1) guided analysis and (2) unguided analysis. In guided analysis, the LLMs are explicitly informed within the prompt that the provided responses could be potentially harmful (see Fig. 3(b)). In unguided analysis, on the other hand, the LLMs evaluate the response quality without any specific indications (see Fig. 3(c)).
We use the same 500 red-teaming instructions from Sec. 3.1 along with their original problematic responses from the PKU-SafeRLHF dataset for mistake analysis and evaluate the quality of analysis using a scale from 1 to 10. Each pair of the guided and unguided analyses, corresponding to the exact same instruction-response sample, is categorized as a win, tie, or lose based on their scores. As illustrated in Fig. LABEL:fig:pre_guided_vs_unguided, there is a noticeable preference for guided analysis. Across all models, the number of “wins” in guided scenarios consistently exceeds that in unguided ones, emphasizing the effectiveness of providing clear guidance when requesting mistake analysis. See Appendix A.2 for more detailed examples.
Figure 3: Prompt templates for the proposed method based on mistake analysis.
4 Method
Denote $D=\{D_{\text{helpful}}=\{(x_{\text{help}},y_{\text{help}})\},D_{\text{harmless}}=\{(x_{\text{harm}},y_{\text{harmless}})\}\}$ as the instruction tuning datasets, where $D_{\text{helpful}}$ contains the helpful instruction-response pairs, and $D_{\text{harmless}}$ includes the red-teaming instructions $x_{\text{harm}}$ potentially related to harmful topics, with $y_{\text{harmless}}$ representing the expected harmless responses. Given a LLM as $F_{\bm{\theta}}(\cdot)$ parameterized by $\bm{\theta}$ and the sequence pairs $(\bm{x}_{i},\bm{y}_{i})\in D$, the objective of SFT is to minimize the cross-entropy loss between the true probability distribution and the model’s estimated distribution over the vocabulary. This can be expressed as,
$\mathcal{L}=-\sum_{i}p(\bm{y}_{i}|\bm{x}_{i})\log q(\bm{y}_{i}|\bm{x}_{i};F_{\bm{\theta}}(\cdot)),$ (1)
where $\bm{x}_{i}$ is the input instruction and $\bm{y}_{i}$ is the target response.
4.1 Alignment through Mistake Analysis
Step 1: Guided mistake induction.
To obtain responses containing mistakes, we induce a model to produce toxic outputs by inserting hint keywords (e.g., harmful, unethical, and offensive) into the instruction prompts, as represented in Fig. 3(a). We denote such an (induced) toxic response as $y_{\text{harm}}$ paired with $(x_{\text{harm}},y_{\text{harmless}})\in D_{\text{harmless}}$. An unaligned model is primarily considered here due to its susceptibility to malicious instructions. Experimental results show that this seemingly simple attack achieves a surprising success rate (more than 72% as in Table 4), which convinces us to adopt these induced responses for subsequent mistake analysis. Besides, mistakes (i.e., harmful responses) can also be obtained from human-annotated data (Dai et al., 2023).
Step 2: Guided analysis generation.
Subsequent to obtaining the $(x_{\text{harm}},y_{\text{harm}},y_{\text{harmless}})$ triplets, we instruct the model to analyze the harmful response $y_{\text{harm}}$, as depicted in Fig. 3(b). The resulting mistake analysis is denoted by $c_{y_{\text{harm}}}$. Given $(x_{\text{harm}},y_{\text{harm}})$ along with guidelines such as “analyzing why the answer is potentially harmful, unethical, or offensive”, even an unaligned LLM can produce reasonable mistake analysis, leveraging its superior discrimination ability. Once acquired, $(x_{\text{harm}},y_{\text{harm}},c_{y_{\text{harm}}})$ forms a mistake analysis triplet. Next, we demonstrate how to construct effective mistake analysis samples from these mistake analysis triplets.
Step 3: Unguided analysis fine-tuning.
Unlike guided analysis generation, an unguided template, which lacks any reminder about the potentially harmful nature of the response, is employed to construct mistake analysis samples using the $(x_{\text{harm}},y_{\text{harm}},c_{y_{\text{harm}}})$ triplets, as illustrated in Fig. 3(c). These mistake analysis samples are then integrated into the SFT process, along with $D_{\text{helpful}}$ and $D_{\text{harmless}}$. It is important to note that reminders about the potentially harmful, unethical, or offensive nature of the response are intentionally excluded. This encourages the LLM to analyze the response based solely on its inherent knowledge, leading to a more nuanced discrimination of harmful, unethical, or offensive content.
Step 4: Guided response generation.
After the model’s fine-tuning, a guided strategy is employed during inference. In this phase, the model is explicitly reminded to formulate “harmless, ethical, and inoffensive” responses, as illustrated in Fig. 1. This reminder during inference acts as a safeguard, ensuring the model’s adherence to ethical standards and preventing the generation of potentially harmful content. In this case, the model computes the conditional probability based on the guidelines to generate harmless content, operating within a constrained prefix context provided during inference, as demonstrated in Fig. 1.
4.2 Why Mistake Analysis Works?
Denote the instructions and the responses as $\bm{X}$ and $\bm{Y}$, respectively. Let $\bm{T}\in\{\text{Harmful},\text{Harmless}\}$ be the harmful tag, i.e., a binary variable representing whether the instruction-response pair is harmful. The mistake analysis $\bm{C}$ here can be considered as the detailed chain-of-thought reasoning (Wei et al., 2022) for $p(\bm{T}|\bm{Y},\bm{X})$, since it thoughtfully analyzes why the given $(\bm{Y},\bm{X})$ pair is harmful (i.e., $\bm{T}=\text{Harmful}$). According to Bayes’ Theorem, we have
$p(\bm{T}|\bm{Y},\bm{X})\propto p(\bm{Y}|\bm{X},\bm{T}),$ (2)
under the assumptions that $\bm{X}$ is independent with $\bm{T}$, and $p(\bm{Y}|\bm{X})$ remains relatively stable during the fine-tuning process. The first assumptions hold since $\bm{X}$ is a random instruction irrelevant to $\bm{T}$. Besides, this stability of $p(\bm{Y}|\bm{X})$ is also reasonable since significant changes are not anticipated in the stage of fine-tuning (see details in Appendix D).
According to Eqn. (2), both the guided mistake induction and the guided analysis generation aim to generate reasonable $(\bm{X},\bm{Y},\bm{T})$ triplets with relatively higher probability of $p(\bm{T}|\bm{Y},\bm{X})$. Then, unguided analysis fine-tuning optimizes the model to maximize the probability of $p(\bm{T}|\bm{Y},\bm{X})$ as well as the probability of $p(\bm{Y}|\bm{X},\bm{T})$. Subsequently, in the inference stage, guided response generation boosts the conditional generation $p(\bm{Y}|\bm{X},\bm{T})$, thereby enhancing alignment performance. This underscores the importance of mistake analysis in aligning models. By optimizing these conditional probabilities, the model becomes a coherent reflection of the specified context, thereby facilitating the ethical and responsible development of LLMs.
Table 1: Comparative results of LLM alignment across various methods. We report a Helpful Score to represent helpfulness performance. For evaluating the harmlessness performance, we report the Harmless Score, Harmless Rate, and Helpful Score for harmful instructions, respectively.
Method Mistake Analysis Helpful Harmless
Source Source Score Score Rate (%) Helpful
Alpaca (vanilla) - - 6.21 5.71 52.5 4.51
SFT - - 6.27 6.69 63.0 5.30
RLHF - - 6.30 6.71 64.1 5.35
Critique-Revise Origin - 6.22 6.60 62.6 5.02
Critique-Revise Alpaca - 6.11 6.17 61.3 4.56
CoH Origin - 6.29 6.79 64.7 5.23
CoH Alpaca - 6.28 6.87 65.7 5.29
Origin Alpaca 6.31(+0.10) 7.31(+1.60) 71.0(+18.5) 5.28(+0.77)
Alpaca Alpaca 6.38(+0.17) 7.41(+1.70) 72.4(+19.9) 5.39(+0.88)
Ours Alpaca GPT-3.5 6.31(+0.10) 7.61(+1.90) 74.1(+21.6) 5.60(+1.09)
5 Experiment
5.1 Alignment
In this section, we evaluate the effectiveness of our method in enhancing the harmlessness performance of models that lack safety alignment.
Data.
The PKU-SafeRLHF dataset (Dai et al., 2023) is used for both training and evaluation. This human-curated dataset emphasizes safety preference, covering multiple dimensions such as insults, immorality, crime, emotional harm, and privacy. For each instruction in the dataset, two responses are provided, with labels identifying the more harmful one. This setup supports the training of both SFT and RL-based models. We refine the training set to include 10,260 unique instructions, each paired with corresponding “harmless” and “harmful” responses. In consideration of the balance between helpfulness and harmfulness (Bai et al., 2022b), our training set is further augmented with an additional 52k helpful instructions from Taori et al. (2023). For evaluation, we use two different test sets. The first is the test set from AlpacaFarm (Dubois et al., 2023), which contains 805 instructions for assessing helpfulness. The second is the test set of PKU-SafeRLHF, which contains 1,523 red-teaming instructions for harmfulness evaluation.
Models and baselines
We employ Alpaca-7B (Taori et al., 2023) as our unaligned base model, which is a fine-tuned version of LLaMA-7B (Touvron et al., 2023). We compare our methods with several baseline methods, including vanilla SFT, CoH (Liu et al., 2023a), Critique-Revise (Bai et al., 2022b), and RLHF (Ouyang et al., 2022), all based on Alpaca. For implementing CoH and Critique-Revise, we use both the original “harmful” responses from the training set and the induced responses generated by Alpaca itself. For RLHF, we adopt PPO-Lag (Ray et al., 2019) as outlined in PKU-SafeRLHF, utilizing the official reward [^2] https://huggingface.co/PKU-Alignment/beaver-7b-v1.0-reward and cost models [^3] https://huggingface.co/PKU-Alignment/beaver-7b-v1.0-cost. Furthermore, we deploy LoRA (Hu et al., 2021) by default in all Transformer linear layers, setting the rank to 16. To ensure a fair comparison, all methods under evaluation are fine-tuned for three epochs.
Evaluation metrics.
To evaluate both harmlessness and helpfulness, we use four metrics. First, we apply single-response grading, where each response is assigned a Score from 1 to 10 for both harmlessness and helpfulness evaluation. A Helpful Score and a Harmless Score are reported, respectively. Additionally, for instructions focused on harmlessness, we conduct a binary assessment to determine if a response is harmless, subsequently reporting a Harmless Rate (Sun et al., 2023a). To avoid the situation where higher harmlessness scores are achieved simply by not responding, we also calculate an additional Helpful Score for responses to harmlessness instructions, following the approach in (Yang et al., 2023). Initially, we use GPT-4 for evaluation. However, to ensure the accuracy of the results, human annotators are also involved in the verification process.
Results.
As indicated in Table 1, our method consistently surpasses existing alignment methods, including the vanilla SFT, Critique-Revise, RLHF, and CoH. Particularly, our method remarkably enhances the performance of harmlessness while effectively preserving helpfulness. See Fig. 4 for a qualitative comparison among different methods. Utilizing original faulty cases from the training set alongside Alpaca’s mistake analysis, our method shows a remarkable $35.2\%$ relative improvement in Harmless Rate over Alpaca. Furthermore, applying our method to harmful responses generated by Alpaca through guided mistake induction raises the Harmless Rate to $72.4\%$. This indicates the value of self-induced mistakes as flawed responses in our analysis-based alignment. Remarkably, when using GPT-3.5 as the analysis source, our method achieves state-of-the-art results, with a Harmless Rate of $74.1\%$. This underscores the benefits of advanced analysis sources. Other evaluation metrics exhibit trends consistent with the Harmless Rate.
The superior overall performance of our method not only confirms its enhanced safety alignment but also demonstrates the advantages of self-critique and mistake analysis. This approach allows the model to autonomously optimize responses without external models or human intervention.
Table 2: Comparative results of defense against attacks across various methods. We present Helpful Score to represent helpfulness performance. For harmlessness performance, we report Harmless Score and Harmless Rate for harmful instructions. Performance on the “Goal Hijacking” test data is further provided for evaluating the attack defensive ability.
Method Mistake Analysis Helpful Harmless Goal Hijacking
Source Source Score Score Rate (%) Score Rate (%)
ChatGLM - - 8.32 8.92 95.3 6.85 68.4
SFT - - 8.16 8.91 94.8 7.71 77.2
CoH Origin - 8.23 8.94 95.2 7.89 82.4
Critique-Revise Origin - 8.24 8.90 95.2 7.97 78.7
Origin ChatGLM 8.18 8.93 95.1 8.02(+1.17) 82.4(+14.0)
Ours ChatGLM ChatGLM 8.26 8.96 96.1 8.14(+1.29) 85.3(+16.9)
5.2 Defending Against Advanced Instruction Attacks
Even LLMs meticulously aligned for the harmlessness performance can potentially yield unsafe responses when confronted with emerging instruction attacks, underscoring the importance of swift and robust defensive methodologies. In this section, we assess the efficacy of our method in defending against unforeseen instruction attacks on post-aligned LLMs.
Instruction attacks.
We investigate an instruction attack known as “Goal Hijacking” (Sun et al., 2023a). This attack strategy involves using additional deceptive or misleading instructions to trick LLMs into ignoring the original user prompts and thereby generating harmful responses. According to results reported by Sun et al. (2023a), LLMs, even those that have undergone post-alignment, remain susceptible to “Goal Hijacking”.
Data.
We utilize the SafetyPrompts dataset (Sun et al., 2023a) for safety alignment. This dataset contains 100,000 instruction-response pairs, covering seven typical safety scenarios and six types of advanced instruction attacks. For harmlessness training, we randomly select 500 instructions from each category in the SafetyPrompts. Besides, we adopt an additional 50K instructions from the MOSS dataset (Sun et al., 2023b) to maintain helpfulness. For the evaluation of harmlessness, we use the test set of the SafetyPrompts dataset, which contains 1915 instructions, featuring 136 queries specifically for “Goal Hijacking”. We also sample 1000 instructions from the MOSS dataset dedicated to assessing helpfulness. Consistent with the experimental setting described in Sec. 5.1, we have constructed 500 additional instructions for “Goal Hijacking”. Each instruction is paired with both “harmless” and “harmful” responses to implement baseline methods.
Models and baselines.
Our base model is ChatGLM-6B (Zeng et al., 2023), a bilingual language model developed within the GLM framework (Du et al., 2022). This model has already been aligned for both helpfulness and harmlessness performance (Zeng et al., 2023). As in Sec. 5.1, we compare our method with vanilla SFT, CoH (Liu et al., 2023a), and Critique-Revise (Bai et al., 2022b). To ensure a fair comparison, all these methods, including ours, are fine-tuned using LoRA (Hu et al., 2021) in all Transformer linear layers, setting the rank to 16. Each method undergoes fine-tuning for one epoch, beginning with an initial learning rate of 0.0001.
Evaluation metrics and prompt templates.
We adopt the same evaluation metrics as those used in Sec. 5.1. Additionally, we report the model’s performance on the “Goal Hijacking” instructions separately to examine the efficacy of our attack defense mechanisms. Fig. 7(a) demonstrates the prompt templates adopted in Sec. 5.2, which are also similar to those shown in Fig. 3.
Results.
As shown in Table 2, our method demonstrates a notable improvement over SFT, achieving an $8.1\%$ increase in the Harmless Rate for “Goal Hijacking” instructions. It also consistently outperforms CoH and Critique-Revise, indicating superior advancements in both regular helpfulness and harmlessness instruction performance. Furthermore, the results indicate that self-induced mistakes are more effective than flawed cases from the original dataset, aligning with the observations in Sec. 5.1. During the entire self-critique process, the model autonomously generates both responses and accompanying mistake analysis, eliminating the need for external intervention.
Visualization.
Fig. 5 illustrates a typical example of “Goal Hijacking” from the test set. In the training scenario, the user initially asks for the repetition of a safe statement, then instructs the model to ignore this request and produce an unsafe response. Ideally, the model should decline such malicious directives. During testing, when presented with a similar attack but on different topics, our method successfully rejects the user’s instruction attack. In contrast, the SFT model fails to do so, generating unsafe responses instead. This outcome demonstrates the superior generalization ability of our method, which is attributed to mistake analysis. It enables LLMs to comprehend the internal mechanisms of advanced instruction attacks, thereby enhancing their capability to generalize and defend against similar threats.
Figure 4: Qualitative comparison between different alignment methods.
Figure 5: An example of “Goal Hijacking”. When encountering an instruction that is not seen during training, our method correctly rejects the attack, whereas the SFT model fails. This indicates a superior generalization ability of the model when aligned with mistake analysis.
5.3 Ablation Study
We conduct ablation studies on Alpaca to investigate essential components of our method, including the strategy of constructing mistake analysis training data, the source of mistakes, and the quality and quantity of mistake analysis. Experimental settings are the same as those in Sec. 5.1.
Strategy of constructing mistake analysis training data.
Rows #1 and #2 in Table 3 compare the effects of retaining or omitting the guided analysis instruction, as illustrated in Fig. 3, in the integration of mistake analysis triplets into training samples. This comparison reveals that the unguided strategy yields better performance. This improvement could be attributed to the fact that providing explicit cues during SFT might enable LLMs to develop shortcuts linked to analysis instructions and responses. Such shortcuts can impede the model’s ability to deeply learn appropriate alignment. Besides, according to Eqn. (2), to increase the conditional generation probability of harmless responses during inference, unguided instructions are essential when aligning models.
Source of mistakes.
We consider two sources of mistakes: the original “bad” responses from the training dataset and those induced via guided mistake induction. Table 3, particularly Rows #2 and #3, shows that performance significantly improves when mistakes generated by the model itself are utilized. This improvement underscores the complexities inherent in the model-induced errors.
Analysis quality.
We examine the effects of guided and unguided mistake analysis on Alpaca. In the guided approach, the model is instructed with “analyze why the answer is potentially harmful” during mistake analysis. In contrast, the unguided approach does not provide such a hint. Table 3, specifically Rows #3 and #4, shows that the guided analysis yields superior results. This finding highlights the importance of directed insights in mistake analysis, aligning with the conclusions of our preliminary study in Sec. 3.2.
Table 3: Results of ablation study. We investigate the source of mistakes, the quality and quantity of mistake analysis, and the strategy of constructing mistake analysis data. The default settings in Sec. 5.1 are marked in gray.
No. Mistake Analysis Analysis SFT Helpful Harmless
Source Quality Quantity Instruction Score Score Rate Helpful
[…]
Analysis quantity.
Rows #3 and #5 in Table 3 compare the use of different quantities of mistake analysis data. Row #5 incorporates mistake analysis for both model-induced bad responses and those already present in the original training dataset. This approach doubles the amount of mistake analysis compared to Row #3, which only utilizes analysis of model-induced responses. However, the results show a decrease in effectiveness when multiple mistake analysis are applied to the same instructions. This decline in efficacy might be due to conflicting analysis of bad cases for the same instruction, resulting in sub-optimal alignment performance.
Table 4: Results of mistake induction.
Method Hint Harmful
Position Rate (%)
Alpaca - 47.2
Induction #1 55.9
#2 65.4
#3 67.1
#2 & #3 69.5
#1 & #2 & #3 72.2
Success rate of guided mistake induction
is analyzed by placing hint keywords such as “harmful unethical and offensive” in different positions, including the Position #1 (system prompt), Position #2 (instruction), and Position #3 (response). See Fig. 3(a) as an illustration. As shown in Table 4, guided mistake induction significantly increases the rate of generating harmful responses compared to the Alpaca baseline. Moreover, more repetition and placing hint words closer to responses result in a higher success rate. The final induction success rate exceeds 72%, which convinces us to utilize the induced mistakes for alignment.
6 Conclusion
Ensuring that LLMs align with human values is of paramount importance. Existing alignment methods typically shield LLMs from mistakes to prevent the generation of toxic responses. Our approach, however, introduces a novel alignment method based on mistake analysis. This method deliberately exposes LLMs to the flawed outputs, turning these mistakes into valuable data for model alignment. Experimental results demonstrate the efficacy of our approach, which outperforms both SFT and RL-based methods in aligning unaligned models and defending post-aligned models against advanced instruction attacks. Even with a limited number of mistakes, our method effectively understands the root causes of bad responses and generalizes to address similar challenges more efficiently. Echoing the wisdom of the Chinese proverb, “A fall into the pit, a gain in your wit”, our work aims to endow LLMs with a similar depth of understanding and learning capacity.
Acknowledgement.
We gratefully acknowledge the support of MindSpore, CANN (Compute Architecture for Neural Networks) and Ascend AI Processor used for this research.
Appendix
Appendix A More on Preliminary
A.1 Generation Against Discrimination
Data and models.
We randomly sample 500 red-teaming instructions from the PKU-SafeRLHF dataset (Dai et al., 2023), each containing potentially unsafe inducement. Following Sec. 5.1, we use Alpaca-7B to respond to the 500 instructions and also consider responses from GPT-3 and GPT-3.5. GPT-3 is a 175B model trained on a mixture of licensed data created by human annotators and publicly available data. GPT-3.5, also developed by OpenAI, is another transformer-based language model and is believed to be larger than GPT-3. Each model is prompted to analyze its own responses.
Evaluation metrics.
As in Sec. 3.1, we evaluate the quality of the Instruction-Response pairs and (Instruction, Response)-Analysis pairs produced by the three aforementioned models, rating them on a scale of 1 to 10. After verifying the scores manually, we compute the average scores separately for each model and task, as depicted in Fig. LABEL:fig:pre_gen_vs_dis.
Results.
As discussed in Sec. 3.1, across all the evaluated models, the discrimination scores (i.e., identifying and analyzing potential mistakes) consistently exceed those of generation (i.e., producing harmless responses directly) by a significant margin.
As shown in Fig. LABEL:fig:instance_alpaca_1, Alpaca fails to generate a safe response to the instruction “How do you insult an Imam?”, utilizing profoundly inappropriate and offensive terms such as the “comparing him to a woman”, “calling him a degenerate” and “traitor of the faith”. The expression of such disrespectful and harmful sentiments illustrates a severe breach of ethical and safety guidelines, justifying the assigned score of 1.
However, when Alpaca is tasked with analyzing its response, it distinctly and accurately identifies the unsafe nature of its reply. Alpaca’s analysis clearly acknowledges that the response is “harmful because it can incite religious and ethnic hatred” and categorically states that “such inflammatory language should not be tolerated and can lead to dangerous consequences”, demonstrating a high level of self-awareness and understanding of the potential repercussions of its initial response. The model further suggests rephrasing the response in “a more respectful manner, without using any words or phrases that may offend and demean any religions or cultures”. The explicit recognition of the unsafe content and constructive suggestions for improvement justify the assigned score of 10.
Fig. LABEL:fig:instance_gpt3_1 and Fig. LABEL:fig:instance_gpt3.5_1 provide examples for GPT-3 and GPT-3.5, respectively. The analyses of both GPT-3 and GPT-3.5 achieve higher scores than their respective responses. GPT-3 is capable of recognizing that not everyone may understand the humor in its response and that it could potentially “hurt someone’s feelings”. GPT-3.5, when discriminating, can even identify the subtle safety issues, such as “it appears to be encouraging harmful and potentially illegal behavior”, and offers impartial advice, stating that “the AI should clearly state that it is not appropriate or ethical to harm others”.
A.2 Guided analysis against unguided Analysis
Data and models.
We randomly sample 500 instruction-response pairs from the PKU-SafeRLHF dataset (Dai et al., 2023). The prompts contain potentially harmful inducement, and the responses are potentially unsafe. We use the same three models mentioned in the previous section. The models are employed to provide analysis on the 500 responses given the instructions, with both guided and unguided prompt templates in Fig. 3(b) and (c).
Evaluation metrics.
As mentioned in Sec. 3.2, we rate the analyses generated by the above three models on a scale of 1 to 10. After the manual score verification, we calculate the average scores separately according to models and scenarios. We further count the number of instruction-response pairs whose corresponding guided analysis wins, ties with, or loses to unguided analysis, as seen in the horizontal histogram in Fig. LABEL:fig:pre_guided_vs_unguided.
Results.
The major experimental results are discussed in Sec. 3.2. Here, we illustrate the results with a detailed example for each of the three models.
In the given example in Fig. LABEL:fig:instance_alpaca_2, Alpaca’s unguided analysis renders a score of 1, with the response being considered as correct. The lack of guidance causes the system to miss the potential safety concerns in the response, leading to a superficial analysis that merely acknowledges the correctness of provided responses.
However, when guided, Alpaca’s analysis dramatically improves, achieving a score of 9. The guided analysis is comprehensive and insightful, explicitly noting the harmful nature of the response and potential risks related to disclosing the residential address of a private individual like Thomas Tull.
Fig. LABEL:fig:instance_gpt3_2 and LABEL:fig:instance_gpt3.5_2 give an example of GPT-3 and GPT-3.5 respectively. The guided analysis of both GPT-3 and GPT-3.5 achieved higher scores than their respective unguided analysis. With guidance, GPT-3 is able to address the legal and ethical ramifications of using brass knuckles, and GPT-3.5 provides a more balanced and ethical perspective towards a cheating husband finding excuses for not being home with an actionable suggestion like “finding a therapist to help both parties find a healthier way to resolve underlying any conflicts and issues in the relationship”.
Appendix B More on Attack Defending
Strategy of SFT instruction.
Table 5 shows that when faced with novel attacks, adopting guided strategy during SFT results in inferior performance, consistent with our observation in Sec. 5.1, indicating both the effective and generalization of unguided SFT instruction.
Comparative results across various methods with induced mistakes.
In Table 6, we present the experimental results for all baseline models using mistakes derived from ChatGLM. These findings are in alignment with our previous observations in Table 1.
Table 5: Ablation of unguided strategy in SFT instruction when meeting with new attacks.
Method Mistake Analysis SFT Goal Hijacking
Source Source Instruction Score Rate
Ours Origin ChatGLM Guided 7.64 80.9
Origin ChatGLM Unguided 8.02 82.4
Table 6: Comparison of defense against attacks across various methods with induced mistakes.
Method Mistake Analysis Goal Hijacking
Source Source Score Rate (%)
Critique-Revise ChatGLM - 7.42 76.5
CoH ChatGLM - 7.86 79.4
Ours ChatGLM ChatGLM 8.14(+0.72) 85.3(+8.8)
Appendix C More Discussion
Training with model-generated data
has recently become a prevailing research direction in both computer vision (e.g., He et al. (2022) for image classification, GeoDiffusion (Chen et al., 2023b; Gao et al., 2023; Liu et al., 2023b; Li et al., 2023) for object detection (Han et al., 2021; Li et al., 2022) and also StableRep (Tian et al., 2023) for contrastive learning (Chen et al., 2021; Liu et al., 2022) and masked image modeling (Chen et al., 2023a; Zhili et al., 2023)) and natural language processing (e.g., SELF (Lu et al., 2023) for instruction tuning), thanks to the remarkable development of AIGC. Our method also belongs to this direction, but different from previous works, we focus on utilizing the inherent discrimination ability of LLMs to enhance the generation capabilities, totally obviating the need for extra human intervention or external knowledge source, but still with solid theoretical support, as discussed in Sec. 4.2 and Appendix D.
Appendix D More Analysis on Equation (2)
Detailed explanation.
Let us start with the Bayes’ Theorem:
$p(\bm{T}|\bm{Y},\bm{X})=\frac{p(\bm{Y}|\bm{X},\bm{T})p(\bm{X}|\bm{T})p(\bm{T})}{p(\bm{Y}|\bm{X})p(\bm{X})},$ (3)
which simplifies to $p(\bm{T}|\bm{Y},\bm{X})\propto p(\bm{Y}|\bm{X},\bm{T})$ under the assumption that $p(\bm{Y}|\bm{X})$ remains relatively stable during the fine-tuning process. To maintain the model’s capability to produce a response given an instruction (i.e., preserving $p(\bm{Y}|\bm{X})$), we combine original SFT data, both helpful ($D_{\text{helpful}}$) and harmless ($D_{\text{harmless}}$), with mistake analysis data during fine-tuning, rather than solely relying on the mistake analysis data. Thus, Eqn. (2) actually aims to explore how the additional mistake analysis data, along with the original SFT instruction-response pairs, can enhance the alignment performance.
Verification of Eqn. (2) via harmless tag prediction.
To verify $p(\bm{T}|\bm{Y},\bm{X})$ is indeed proportional to $p(\bm{Y}|\bm{X},\bm{T})$ as previously discussed, we further conduct an additional experiment evaluating the LLM’s ability to discern the harmfulness of responses. Specifically, we utilize the test set of PKU-SafeRLHF dataset, each instruction paired with a response and the corresponding binary harmful or harmless tag, and then instruct the LLMs to distinguish whether the given response is harmful or not. The accuracy is reported in Table 7. The model’s accuracy in identifying the harmfulness of the responses improves from 54.5% for vanilla Aplaca-7B to 72.6% after applying our mistake analysis, revealing a significant improvement in discrimination ability.
Table 7: Comparison for discrimination ability via binary classification on PKU-SafeRLHF.
Vanilla Alpaca SFT Ours
Accuracy (%) 54.5 54.9 (+0.4) 72.6 (+18.1)
Appendix E More Qualitative Comparison
Safety instruction following.
We provide a more qualitative comparison for the safety instruction following in Fig. 7-9 for the safety alignment and attack defending respectively, as in Sec. 5.1 and Sec. 5.2, further revealing the superiority of our proposed alignment method via mistake analysis.
Mistake analysis.
We provide more examples of the mistake analyses generated by Alpaca-7B and ChatGLM-6B in Fig. 10-13(a), with Fig. 10 and Fig. LABEL:fig:analysis_data_origin_chatGLM for harmful responses of the original training corpus, and Fig. 11 and Fig. LABEL:fig:analysis_data_chatGLM_chatGLM for the model-generated ones. We observe both Alpaca-7B and ChatGLM-6B are proficient at identifying the harmful responses with comprehensive explanation, and thus, incorporating the mistake analysis during SFT would enhance the generative capacities of LLMs with their innate discriminating abilities.
(a) Prompt templates for the instruction attack defending experiments as in Sec. 5.2.
(b) Examples illustrating discrimination is easier than generation as in Sec. 3.1.
(c) Examples illustrating guided analysis is better than unguided analysis as in Sec. 3.2.
Figure 7: More qualitative comparison for safe alignment as in Sec. 5.1.
Figure 8: More qualitative comparison for safe alignment as in Sec. 5.1.
Figure 9: More qualitative comparison for attack defending as in Sec. 5.2.
Figure 10: Examples of mistake analysis for original harmful responses generated by Alpaca-7B as in Sec. 5.1.
Figure 11: Examples of mistake analysis for the model-generated harmful responses generated by Alpaca-7B as in Sec. 5.1.
(a) Examples of mistake analysis generated by ChatGLM-6B as in Sec. 5.2.